If I buy a put option with a price of $1, with a strike price of $110, and the stock price ends up being $90, what is my holding period return?

How would this return be different if a dividend of 2.5% was received?

Subject: Portfolio management
Answer #1

A put option is a contract in which the buyer has the right to sell the stock at a specified price.

Since the current price is $90 and the user has put option, hence he can sell the stock at $110. Thus, it is beneficial for him to use this put option.

Now assuming here, that the user decides to use this put option, following will be his holding period return (HPR)

HPR (per stock) = ( Selling price - Buying price - price paid to buy put option )/ buying price x 100

HPR = (110-90-1)/90 *100

HPR= 19/90*100 = 21.11%

Now this return would be different if the user also receives dividend of 2.5%

Dividend received = 2.5% *90 = $2.25

HPR (per stock) = ( Selling price +dividend received - Buying price - price paid to buy put option )/ buying price x 100

HPR = (110+2.25-90-1)/90 *100

HPR= 21.25/90*100 = 23.61%